the length of a rectangular pool is 4 times its width. If the perimeter of the pool is 140 meters, find the length and width of the pool.
Answer by swincher4391(1107)

L = 4W
2L + 2W = 140
2(4W) + 2W = 140
8W + 2W = 140
10W = 140
W = 14
L = 56
56x14